# Coherent Control of an Optical Quantum Dot Using Phonons and Photons

### v1

@@ 1 System Hamiltonian
Figure S1: (a) Calculated TLS occupancies for five different values of $\phi$ (solid curves). The $g=0$ trajectory (which does not depend on $\phi$) is also shown for reference (dotted curve) (b) Calculated TLS occupancies for $\phi$=0 (blue curve) and for $\phi$ averaged over eight uniformly-spaced values between 0 and $2\pi$ (orange curve). All calculations in panels a,b were performed for an idealized square pulse with the following parameters: $g/2\pi$=1 GHz, $\Omega_{0}$=1 GHz. (c,d) Occupancies and occupancy ratios ($c_{g}(t)$) calculated for the same set of parameters as in manuscript Fig. 3d for (c) $\phi$=0 and for (d) $\phi$ averaged over eight values between 0 and $2\pi$.
@@ 2 Experimental Setup
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. Nonresonant and resonant pump beams were shaped outside the cryostat using single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with an approximately 10 kHz linewidth at 970 nm. The nonresonant pump was a 632 nm diode laser. Pump beams illuminate the QD through the objective. QD luminescence was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant pump) or polarization filtering (for measurements involving a resonant pump) with at least 106 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant for all measurements except for Fig. 2c of the manuscript.
@@ 2.1 Bias-dependent QD Photoluminescence Spectra
Figure S2: QD PL spectrum with varying Schottky bias voltage. Two to three QDs contribute to this spectrum. The QD investigated in the manuscript is indicated by a white arrow.
@@ 2.2 Continuous-wave Resonance Fluorescence Spectra
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. manuscript Fig. 2a) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence counts at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 2 nW at the sample was used, corresponding to approximately 1/10 the measured saturation level of the QD. For manuscript Fig. 2d (upper panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.518 GHz and external microwave driving powers between -50 dBm and -35 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit for each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration data are shown in Fig. S3.
For manuscript Fig. 2d (lower panel), the “resonant” pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 5 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. In manuscript Fig. 2d (lower panel), such spectra were again measured without SAW driving and with SAW driving at 3.518 GHz and the same powers as in manuscript Fig. 2d (upper panel).
@@ 2.3 Pulsed Optical Measurements and Time-domain QD Dynamics
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “exponential” decaying pulses, optical pulses were generated by passing a CW “resonant” optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG). The rise-time, fall-time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be better than $10^{4}$. The frequency bandwidth of the output pulses was measured to be 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). Measured pulse profiles shown in the manuscript may differ slightly from the pulse experienced by the QD due to small system reflections outside the cryostat. The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. We typically used 30-60 second total integration times, over which we collected typically several million total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 1500 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations.
The time-resolved QD dynamics shown in the manuscript are presented after subtracting the experimentally measured pump contribution. This subtraction process is illustrated in Fig. S4, where black curves represent the directly measured data including pump and QD luminescence contributions, gray regions represent the measured pump contribution, and blue curves represent the difference which we associated with the QD contribution.
Figure S4: A series of raw TCSPC measurements showing measured pump background (gray filled region), QD signal (black curve) and pump-subtracted QD signal (blue curve). Each panel includes a different set of experimental parameters specified above each panel.
@@ 3 Estimating QD Occupancies from Experimental Time-Domain Data
QD occupancies were derived by comparing count rates from time-resolved QD luminescence data to a reference count rate. The reference count rate was derived by performing a set of time-resolved QD luminescence measurements with 130 ps optical pulses and widely varying optical pump powers and analyzing the resulting Rabi oscillations in the signal. Experimental Rabi oscillations were compared to numerical calculations of the QD occupancy under a 130-ps square pulse. The calculation includes dephasing assuming that the QD dephasing is purely radiative. (The measured 1 ns radiative lifetime corresponds to a 321 MHz decay rate.) This calibration is illustrated in Fig. S5. Note that experimental occupancies will be overestimated if the QD experiences additional dephasing. Nonetheless, this scaling is irrelevant for the enhancements presented in the manuscript.
Figure S5: Left panel: Raw TCSPC data acquired with $\Delta=0$ and a 130 ps pulse duration. “Maximum counts” refers to the counts recorded in a 16-ps time bin at 140 ps, i.e., immediately after the pump pulse. Right panel: Counts recorded per second within a 16-ps time bin at 140 ps as a function of average optical power (open black square markers). Single-time-bin counts are converted to an expected QD occupancy $\langle\rho_{ee}\rangle$ by comparing the resulting Rabi oscillations to theoretical calculations including lifetime-limited dephasing at rate $\gamma_{\text{QD}}/2\pi$=320 MHz (solid black curve).

### v2

@@ 1.3 Mechanical phase
Figure S1: (a) Calculated TLS occupancies for five different values of $\phi$ (solid curves). The $g=0$ trajectory (which does not depend on $\phi$) is also shown for reference (dotted curve) (b) Calculated TLS occupancies for $\phi$=0 (blue curve) and for $\phi$ averaged over eight uniformly-spaced values between 0 and $2\pi$ (orange curve). All calculations in panels a,b were performed for an idealized square pulse with the following parameters: $g/2\pi$=1 GHz, $\Omega_{0}$=1 GHz. (c,d) Occupancies and occupancy ratios ($s_{g}(t)/s_{g=0}(t)$) calculated for the same set of parameters as in manuscript Fig. 3d for (c) $\phi$=0 and for (d) $\phi$ averaged over eight values between 0 and $2\pi$.
2 Sample Fabrication
Our devices are fabricated on InAs quantum dot (QD) samples grown by molecular beam epitaxy (MBE) on a commercial GaAs wafer. The vertical structure consists essentially of a p-i-n diode above a distributed Bragg reflector (DBR) comprising 22 pairs of alternating AlAs/GaAs layers. Self-assembled InAs QDs are grown in the intrinsic (i) region between the highly doped p and n GaAs regions. The n (GaAs:Si at approximately $2\times 10^{18}$ cm-3) layer is 46.8 nm thick at a depth 300 nm below the surface. The InAs quantum dots (QDs) are grown 25 nm above this n layer. The p (GaAs:C at approximately $2.5\times 10^{19}$ cm-3) is 45 nm thick and lies at the sample surface. A 220 nm thick AlAs/GaAs digital alloy (current blocking layer) is grown between QDs and the p layer.
The MBE-grown wafer is then processed in a cleanroom for QD p-i-n gate control and active surface acoustic wave (SAW) cavities. SAW mirror trenches are defined by electron beam lithography (EBL) and etched by reactive-ion etching (RIE). Large wirebond pads are defined for electrically contacting the n and p layers. Ohmic contacts to the n layer are made by etching vias around the SAW resonators down to approximately 80 nm above the n layer, and then depositing approximately 400 nm of AuGe/Ni/Au by electron-beam evaporation. Ohmic contacts to the p layer are made by depositing approximately 200 nm of Pt/Au to the p surface. The metals are thermally annealed in forming gas (95% Ar and 5% H2) at 430 C for 1 minute. The n regions around each SAW device are selectively removed by completely etching the surface to a depth 375 nm (25 nm below the n layer) so that this conductive layer does not interfere with on-chip microwave electronics. The n layer is retained only under SAW cavities (including mirrors and interdigital transducers, IDTs) and regions where the n contact metals were previously deposited. The coplanar waveguide (CPW) geometry is then defined by photolithography and formed by depositing 400 nm of Ti/Au followed by a lift-off process. The thickness of the CPW metal is chosen to be thicker than the etched depth so that it is connected across the n etch step. Finally, 20 nm of Al is deposited to make the IDTs. The IDT structure is defined by EBL and then formed using a lift off process.
@@ 3 Experimental Setup
All of our QD optical measurements were performed using a 5 K optical cryostat with an internal cryogenic 0.81 NA objective. Nonresonant and resonant pump beams were shaped outside the cryostat using single-mode optical fiber components and free-space optical components. The resonant pump beam was a tunable external-cavity diode laser with an approximately 10 kHz linewidth at 960 nm. The nonresonant pump was a 632 nm diode laser. Pump beams illuminate the QD through the objective. QD luminescence was collected by the same objective and isolated from the pump predominantly by spectral filtering (for measurements involving a nonresonant pump) or polarization filtering (for measurements involving a resonant pump) with greater than 107 pump rejection. Luminescence was collected into single-mode optical fiber and various frequency- and time-domain measurements were performed using procedures detailed below. The QD bias voltage was held constant at 0.41 V for all measurements presented in the manuscript.
@@ 3.1 Bias-dependent QD Photoluminescence Spectra
Figure S2: QD PL spectrum with varying p-i-n bias voltage. Approximately three QDs contribute to this spectrum. The brightest three lines are associated with three distinct charge states of the single QD investigated in this work: $X^{0}$, neutral exciton; $X^{+1}$, single positively charged exciton; $X^{-1}$, single negatively charged exciton. The specific charge state investigated in this work is indicated by the white arrow.
@@ 3.2 Continuous-wave Resonance Fluorescence Spectra
Continuous-wave (CW) resonance fluorescence (RF) spectra (e.g. manuscript Fig. 2d) were recorded by slowly sweeping the frequency of a tightly focused resonant pump beam over the QD’s exciton resonance while holding the QD bias constant. Luminescence counts at each frequency were recorded using a superconducting nanowire single photon detector (SNSPD) mounted in a dilution refrigerator. A pump power of approximately 500 pW at the sample was used, corresponding to approximately 1/10 the saturation level of the QD, providing a resonance fluorescence signal with approximately 150 kcounts/second when pumping the QD resonantly. For manuscript Fig. 2d (upper panel), such spectra were recorded without SAW driving and with SAW driving using an external coherent microwave source at approximately 3.5881 GHz and external microwave driving powers of approximately -41.5 dBm and -36.5 dBm. External microwave driving powers were converted to coupling rates $g$ by a series of calibration data where the modulation index $\chi$ was fit from optical spectra at each microwave driving power and $g$ was derived by the relationship $g$=$\omega_{\text{SAW}}\chi$. The calibration results are shown in Fig. S3.
For manuscript Fig. 2d (lower panel), the “resonant” pump was held at a constant frequency detuning of $\Delta$=$-\omega_{\text{SAW}}$. A pump power of approximately 2 nW was used. The frequency spectrum of scattered photons was measured by passing the collection port through a tunable Fabry-Perot etalon filter with a 600 MHz linewidth before being counted on an SNSPD. Such spectra were again measured without SAW driving and with SAW driving at 3.5881 GHz and the same microwave drive powers as in manuscript Fig. 2d (upper panel).
@@ 3.3 Pulsed Optical Measurements and Time-domain QD Dynamics
The time evolution of the QD dynamics was measured using standard time-correlated single-photon counting (TCSPC) techniques with a two-channel picosecond event timer and TCSPC module. For “gradually decreasing” pulses, optical pulses were generated by passing a CW optical pump (same as described above) through two cascaded electro-optic modulators (EOMs). The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). The EOMs were driven by amplified pulses from an approximately 100-ps square-wave electronic pulse generator (EPG). The EPG was externally triggered using an arbitrary waveform generator (AWG). The rise-time, fall-time and width of the output pulses were measured to be approximately 15 ps, 15 ps, and 130 ps respectively. The on/off contrast was measured to be about $10^{5}$. The frequency bandwidth of the output pulses was measured to be approximately 7 GHz. To decrease the bandwidth (and increase the pulse duration), the pulses were transmitted through a 600 MHz tunable Fabry-Perot etalon which was tuned for maximum transmission at the pulse’s center frequency. The resulting temporal envelope is shown in Fig. 3a (top panel). For square pulses, the EOMs were directly driven using a high-speed AWG (64 GS/sec) with approximately 30 ps rise time and pulses of arbitrary length. The pulse center frequency was detuned by $\Delta$ with respect to the QD’s resonance frequency (detuning specified in each measurement). Measured pulse profiles shown in the manuscript may differ slightly from the pulse experienced by the QD due to small system reflections inside and outside the cryostat. The time distribution of the QD luminescence was measured by again detecting the luminescence with an SNSPD and time-correlating the counts on the TCSPC module with respect to a trigger signal from the AWG. In both pulse cases, the repetition rate was 34.085 MHz. We typically used 30-60 second total integration times, over which we collected typically tens of millions total photons. This type of measurement was performed with a wide range of average optical input powers (between approximately 0.25 nW to 300 nW). Pump powers were converted to resonant Rabi rates by looking at the resulting time-domain oscillations with zero detuning.
The time-resolved QD dynamics shown in the manuscript are presented after subtracting the experimentally measured pump contribution. Nonetheless, in all cases, unrejected pump contributions were $\lesssim$1:100 compared to the measured signal and negligibly impact our results.
@@ 4 Estimating QD Occupancies from Experimental Time-Domain Data
QD occupancies were derived by comparing count rates from time-resolved QD luminescence data to a reference count rate. The reference count rate was derived by performing a set of time-resolved QD luminescence measurements with 130 ps optical pulses and widely varying optical pump powers and analyzing the resulting Rabi oscillations in the signal. Experimental Rabi oscillations were compared to numerical calculations of the QD occupancy under a 130-ps square pulse. The calculation assumes that the QD experiences no additional dephasing beyond the radiative lifetime limit. (The measured 1 ns radiative lifetime corresponds to a 321 MHz decay rate.) This calibration is illustrated in Fig. S4. Note that experimental occupancies reported here will be overestimated if the QD experiences additional dephasing. Nonetheless, excellent contrast observed in measured power-dependent Rabi oscillations (Fig. S4) indicate that additional dephasing beyond the radiative lifetime limit may be very small for this particular QD. This scaling is, however, irrelevant for the enhancements presented in the manuscript.
Figure S4: Left panel: Raw TCSPC data (blue curve; units counts/second per time bin) acquired with $\Delta=0$ and a 130 ps pulse duration (gray region) and 16 ps time bin. “Maximum counts” refers to the counts recorded in a 16-ps time bin at 140 ps, i.e., immediately after the pump pulse. Right panel: Counts recorded per second within a 16-ps time bin at 140 ps as a function of average optical power (open black square markers). Single-time-bin counts are converted to an expected QD occupancy $\langle\rho_{ee}\rangle$ by comparing the resulting Rabi oscillations to theoretical calculations including lifetime-limited dephasing at rate $\gamma_{\text{QD}}/2\pi$=320 MHz (solid black curve).
Figure S5: Data and calculations shown in Fig. 3 of the manuscript presented in 2D.
Figure S6: Numerically calculated phonon-assisted QD excitation dynamics without additional dephasing and no spectral drift. I.e., the QD dephasing is assumed to be radiative-lifetime limited with $\gamma/2\pi$=320 MHz. All other conditions are identical to those used for manuscript Figs. 3c,d. Left column: gradually decreasing pulse. Right column: square pulse. The “enhancements” shown for the gradually decreasing pulse shape are presented in a log scale since maxima vary by several orders of magnitude over the pulse duration.
Figure S7: Scattering spectra from the same single QD excited with a continuous-wave narrowband red-detuned pump at $\Delta$=$-\omega_{\text{SAW}}$=--3.581 GHz. Spectra were measured with a frequency-tunable spectral filter with a 25 MHz transmission bandwidth. Gray curve: no SAW drive. Orange curve: SAW drive with $g$$\approx$1 GHz. The narrow “coherent” Rayleigh scattering feature is seen at $\Delta$=--3.581 GHz in both spectra. When the SAW drive is on, some of this “coherent” scattering component is frequency up-converted by phonons to leading to a narrow spike in the frequency at $\Delta$=0. This process also leads to a finite average QD occupancy and thus a broader luminescence feature with a width of approximately 1 GHz around $\Delta$=0. We suspect the small “dips” between the narrow and broad emission features at $\Delta$=0 originate from phase interference between two scattering channels, but the origin is still under investigation.